OpenStudy (anonymous):

Help me simply this?? x^3-125

OpenStudy (anonymous):

Use the fact that x^3 + y^3 = (x + y)*(x^2+y^2-xy), so x^3 - 125 = x^3 + (-5)^3 = (x-5)*(x^2-5x+25).

OpenStudy (anonymous):

Where did the y come from?

OpenStudy (anonymous):

i thinks it is o

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

Here are the formulas you will need to use: x^3 - y^3 = (x-y)(x^2 +xy +y^2) x^3 + y^3 = (x+y)(x^2 -xy +y^2) 1) x^3 - 125 is a difference of two cubes. x^3 - 125 = x^3 - 5^3 = (x-5)(x^2 + 5x + 25) 2) x^3 + 8y^3 is a sum of two cubes x^3 + 8y^3 = x^3 + (2y)^3 = (x+2y)(x^2 - 2xy + 4y^2)

OpenStudy (anonymous):

You just copy and pasted that from yahoo answers....

OpenStudy (anonymous):

I don't understand where the 5^3 comes from... There has to be an easier way to do this rather than using a formula...

OpenStudy (the_fizicx99):

There isn't. :/ Use wolfram alpha if it helps

OpenStudy (anonymous):

Then there must be an easier way to explain it?

OpenStudy (anonymous):

In general- Difference of Cubes: (a^3-b^3)= (a-b)(a^2+ab+b^2) Sum of Cubes: (a^3+b^3)= (a-b)(a^2-ab+b^2) 1) Using the difference of cubes, (a=x, b = 5) we get (x-5)(x^2+5x+25) 2. Using the sum of cubes, (a=x, b = 2y) we get (x+2y)(x - 2xy + 4y^2) Hope this helps!

OpenStudy (anonymous):

Dude just quit copy and pasting yahoo answer crap, I've already been to all of those

OpenStudy (haseeb96):

x^3-125=0 x^3=125 x^3=5^3  this is the correct answer